12 Which of the following mixtures produces a buffer solution?
0.300 M NaOH + 0.100 M HCl (50 mL each)
0.300 M $KNO_3$ + 0.100 M NaCl (50 mL each)
0.300 M $NH_4Cl$ + 0.100 M HCl (50 mL each)
0.400 M $CH_3COOH$ + 0.100 M $CH_3COONa$ (50 mL each)



13 Buffering is:
due to LeChatelier's Principle
is a process that is done by a solution made up of a weak acid and its salt
influenced by the common ion effect
all of these



14 What’s the concentration of Ag+ ion in a saturated silver chloride
solution? $K_{sp}$ = 1.56 x$10^{-10}$.
0.0000125 M
4.90 M
0.000339 M
none of these



15 How many coulombs of charge pass through a cell if 2.40 amperes of
current are passed through the cell for 85.0 minutes?
204 C
0.133 C
12200 C
2120 C
1
Expert's answer
2016-05-31T12:17:03-0400
12. 0.400 M CH3COOH + 0.100 M CH3COONa (50 mL each)
According to definition, buffer solution is the mixture of the salt of weak acid (or base) with this acid (or base).
13. all of these.
14. The concentration of silver ions is:
K_sp=[Ag^+ ][Cl^- ]=[Ag^+ ]^2
[Ag^+ ]= √(2&K_sp )=1.25 ∙10^(-5) M
(0.0000125 M)
15. q = I*t = 2.40 (A) 85 (min) 60 (sec/min) = 12240 C. (12200 C)